# Speeding up an unsteady flow simulationM. HANEK, J. PAPEŽ, and J. ŠÍSTEK

### v1

@@ 7 Conclusions
We have focused on numerical techniques for the solution of time-dependent problems of incompressible flows using time stepping. In particular, we have studied the benchmark problem of flow around a unit sphere with Reynolds numbers 100 and 300. These Reynolds numbers result in fundamentally different regimes with respect to the dependence on time. While for Re = 100, the solution exhibits a transient behaviour converging to a steady solution, the solution for Re = 300 loses stability and becomes periodic in time.
@@
We have applied the incremental pressure-correction method which leads to solving five systems of linear equations in each time step. Out of these, we have focused on the solution of the Poisson problem for pressure corrector which is the most challenging part of the process. From the linear algebraic viewpoint, the problem leads to solving a sequence of algebraic systems with a constant symmetric positive-definite matrix and a right-hand side vector that changes in each time step. Importantly, in the transient regime for Re = 100, the systems within the sequence are getting closer to each other, while this property does not hold for the periodic regime with Re = 300.
The systems are solved in parallel using the nonoverlapping domain decomposition method. In particular, we have used the three-level BDDC preconditioner within PCG due to its scalability to very large problems and parallel supercomputers. The standard setting of the three-level BDDC preconditioner provided a baseline for our study, the aim of which has been accelerating the computation across the time steps. Three main algorithmic components have been tested, and their effects on the time-to-solution have been evaluated. First, we explored the effect of initial guess and the stopping criterion. It has been shown that in order to make use of a good initial approximation of the solution, the stopping criterion has to be modified from the usual one based on the initial residual to be independent of the initial guess. In our approach, we have used a stopping criterion normalized by the norm of the right-hand side vector. With this criterion, the Krylov subspace method is capable of decreasing the number of iterations with improving the initial guess, such as using the solution from the previous time step. This capability of the solver turned out to be crucial for saving more than 50 percent of computational time in the transient regime for Re = 100.
The second component of the study has been the deflation within PCG combined with Krylov subspace recycling. We have tested four strategies of constructing the deflation basis. Two strategies were based on storing previous search directions, while the other two were based on approximating eigenvectors of the preconditioned system matrix by harmonic Ritz vectors. However, there is a significant difference in our approach from the standard use of the latter. Namely, with a strong preconditioner such as BDDC, the standard approach for deflating the smallest eigenvalues need not work well, and we propose constructing the deflation basis from approximations of the eigenvectors corresponding to the largest eigenvalues. In this regime, the deflation acts as another level of preconditioning, further lowering the upper part of the spectrum of the preconditioned system matrix. The increased cost of each iteration due to the deflation has been compensated by the reduced number of iterations needed to converge.
The last component towards accelerating the method has been the adaptive selection of the coarse space in BDDC. Adaptive BDDC allows to adjust the strength of the preconditioner. Based on solving a number of eigenvalue problems for each pair of subdomains sharing a face, the method constructs a coarse space so that the resulting condition number is approximately bounded by the threshold $\tau$. The smaller the threshold, the more coarse degrees of freedom are defined leading to a smaller number of iterations. However, the cost of solving the coarse problem increases making each iteration more expensive. Due to a significant time spent in the setup of the adaptive BDDC preconditioner, the approach is beneficial for longer runs.
@@
We have shown the synergy of the three approaches for the acceleration of the time-dependent solution. In particular, using the appropriate stopping criteria, previous solution as the initial guess, deflation in the PCG and the adaptive coarse space construction in the multilevel BDDC preconditioner, we have achieved a significant reduction of the computational time. For the transient regime, the reduction was more than one half solely by using the appropriate stopping criterion and more for the other two techniques. For the more challenging problem with the periodic solution, we have saved about one quarter of the computational time for the smaller mesh with 1.4M unknowns for pressure and about one half of the computational time for the larger problem with 15.6M unknowns. Hence, the advanced techniques of Krylov subspace recycling and adaptive coarse space construction seem to be even more beneficial for more challenging problems which require more PCG iterations for each time step.
Finally, let us mention that there is an optimal size of the deflation basis in the deflated PCG. A similar effect holds for threshold $\tau$ within the adaptive BDDC preconditioner. While we have approximately identified these optima in our computations, the results show that these optima are not sharp, and comparable speed-ups can be achieved for a range of values of the parameters. In this sense, the proposed approach seems to be quite robust.

### v3

@@ 7 Conclusions
We have focused on speeding up the solution of the Poisson problem for pressure corrector in time-dependent incompressible flows problems. From the linear algebraic viewpoint, the problem leads to solving a sequence of algebraic systems with a constant symmetric positive-definite matrix and a right-hand side vector that changes in each time step. A baseline for our study was provided by the three-level BDDC preconditioner within PCG. Three algorithmic components have been tested, and their effects on the time-to-solution have been evaluated.
We first studied the influence of the initial guess and the stopping criterion. To benefit from a good initial approximation, the stopping criterion must be independent of the initial guess. We therefore used a criterion normalized by the norm of the right-hand side. This allows the Krylov method to reduce the iteration count as the initial guess improves (e.g., when reusing the previous time-step solution), resulting in more than 50% savings in computational time in the transient regime for Reynolds number Re = 100.
The second part of the study focused on deflation in PCG combined with Krylov subspace recycling. We tested four strategies of constructing the deflation basis: two based on storing previous search directions and two based on harmonic Ritz approximations of eigenvectors. Unlike the standard approach, which targets the smallest eigenvalues, we exploit the fact that preconditioners such as multilevel BDDC yield spectra bounded below by one and therefore we propose constructing the deflation basis from approximations of eigenvectors associated with the largest eigenvalues. This shifts down the upper end of the spectrum and, despite the higher per-iteration cost, reduces the total number of iterations and overall computational effort.
The final acceleration component was the adaptive selection of the BDDC coarse space. Adaptive BDDC tunes the preconditioner strength by solving local eigenvalue problems on subdomain interfaces and constructing a coarse space that reduces the condition number according to a threshold $\tau$. Smaller values of $\tau$ introduce more coarse degrees of freedom, decreasing the iteration count but increasing the per-iteration cost. Because the setup of the adaptive BDDC preconditioner is relatively expensive, this approach is particularly advantageous for longer simulations.
@@
We demonstrated the synergy of three acceleration techniques for time-dependent problems. Using an appropriate stopping criterion, the previous time-step solution as the initial guess, deflation in PCG, and adaptive coarse space construction in the multilevel BDDC preconditioner led to substantial reductions in computational time. In the transient regime (Reynolds number Re = 100), more than 50% of the time was saved by the stopping criterion alone. For the periodic regime with Re = 300, we achieved savings of about 25% for the smaller problem with 1.4M pressure unknowns and over 40% for the larger problem with 15.6M unknowns. These results indicate that Krylov subspace recycling and adaptive coarse spaces become increasingly beneficial for more challenging problems requiring many PCG iterations. While the individual components are well known, their synergistic combination has not been previously reported in the literature.
Finally, we note that both the size of the deflation basis in deflated PCG and the threshold $\tau$ in adaptive BDDC admit optimal values. Please note that the goal of the paper was not to show the optimal setup of the parameters. Instead, we identified them approximately and showed that they are not sharp, as comparable speed-ups are obtained over a range of parameter values. This indicates that the proposed approach is robust and broadly applicable to simulations of unsteady incompressible flows.
@@ Acknowledgements
Data availability
Data will be made available on request.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.